If you travel for 2 hours at a rate of 15 mph, and you travel for 6 hours at a rate of 5 mph, what is your average speed and how do you solve it?
Answer by stanbon(75887)

If you travel for 2 hours at a rate of 15 mph, and you travel for 6 hours at a rate of 5 mph, what is your average speed and how do you solve it?
--------------
average speed = (total distance)/(total time)
-------
average = (2*15+6*5)/(2+6) = 60/8 = 7.5 mph
========================
Cheers,
Stan H.